Determine the area enclosed by y =x and

y =x^2


1
Expert's answer
2021-07-06T10:52:04-0400

"y=x.... (1)"

"y=x^2.....(2)"

From equation (1) and (2)

"x=x^2"

or "x-x^2=0"

"x=0,1"

But "y=x"

"y=0,1"

"(0,0)" and "(1, 1)" are points of intersection

"A=\int_0^1(x-x^2) dx"

"=[{x^2\/2} - {x^3\/3}]_0^1"

"=[1\/2-1\/3]-[0-0]"

"=1\/6" square units